OpenStudy (anonymous):

find the closest point of curve: x^4+y^4=0 to the line : x+8y = 24.

OpenStudy (anonymous):

x^4 + y^4 = 0, if x = 1, what is y? If x= -1, what is y? If x= 0, what is y?

OpenStudy (anonymous):

:-O I didn't understand

OpenStudy (anonymous):

You have an equation x^4 + y^4 = 0 and I am asking you what is the value of y when x= -1, 0 and 1.

OpenStudy (anonymous):

for x=-1 y is 1 for 1 y is -1 and for 0 y is 0

OpenStudy (anonymous):

You say that when x is 1, y is -1 1^4 = 1, what does (-1)^4 equal?

OpenStudy (anonymous):

-1*-1*-1*-1 = ?

OpenStudy (anonymous):

You do know that when you multiply a minus number by a minus number you get a positive number?

OpenStudy (anonymous):

Oh, you deleted your post, OK.

OpenStudy (anonymous):

so how can these answers be related to my question?

OpenStudy (anonymous):

When you answer my question,I will tell you.

OpenStudy (anonymous):

OK, I guess you are unable to multiply -1 by itself 4 times, the answer is 1. But 1+1 is not equal to zero. And if I put x = -1, same thing. So the only solution to your first equation (which you called a curve) is (0,0).

OpenStudy (anonymous):

why did u try just these 3 integers?

OpenStudy (anonymous):

I could have put a, -a and 0, I chose them as being representatives of all possible values of x.

OpenStudy (anonymous):

what about my next equation that is a line, u didn't even use it

OpenStudy (anonymous):

Whatever you next equation, there is only one solution to the first equation, so that must be the nearest point (it is the ONLY point).

OpenStudy (anonymous):

but this point (0,0) in not on this line because 0+0 != 24 am I right?

OpenStudy (anonymous):

Correct, you asked for the closest point.

OpenStudy (anonymous):

but it may be nearest point yeah thanks

OpenStudy (anonymous):

If your second equation was a line passing through the origin, the point would be coincident.

OpenStudy (anonymous):

so as last question, what do u think about the figure of first equation? u said just one point is correct on this equation so is that a point??

OpenStudy (anonymous):

I said before, (0,0) (which is a point like any other although it is also the origin).